Product of the same integer with itself
Answer #1

The product depends on the sign of the integer. For suppose assume two integer a and b, their magnitude is equal |a| = |b|.i.e, for example a = 10 then also b=10.

Now considering the sign of the two integers,

(+a) ×(+b) =ab,

(+a) ×(-b) = -ab,

(-a) ×(+b) = -ab,

(-a) ×(-b) = ab.